Find the surface area, to the nearest square inch, of a cone with radius 6 in. and slant height 8 in.
Answer by palanisamy(496)

Given, the radius of the cone r = 6 in
the slant height l = 8 in
The curved surface = pi*r*l
=(22/7)*6*8
= 150.86 sq in
Base area = pi*r^2
=(22/7)*6^2
= 113.14
Therefore total surface of the cone = 150.86+113.14
= 264 sq in